OpenStudy (anonymous):

Find the vertex of the parabola whose equation is y = x^2 + 6x + 2. What is the line of symmetry for the parabola whose equation is y = 3x^2 + 24x - 1? What is the line of symmetry for the parabola whose equation is y = x^2 + 10x + 25? Find the vertex of the parabola whose equation is y = -2x^2 + 8x - 5. What is the line of symmetry for the parabola whose equation is y = x^2 - 12x + 7? Find the vertex of the parabola whose equation is y = x^2 - 4x + 6.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

What is the line of symmetry for the parabola whose equation is y = x^2 + 10x + 25?

OpenStudy (anonymous):

\(-\frac{b}{2a}=-\frac{10}{2\times 1}=-5\) is what you are looking for

OpenStudy (anonymous):

line is \(x=-5\)

OpenStudy (anonymous):

we havent done the first one yet silly! ;] @satellite73

OpenStudy (anonymous):

Find the vertex of the parabola whose equation is \(y = -2x^2 + 8x - 5\) first coordinate of the vertex is \(-\frac{b}{2a}=-\frac{8}{2\times (-2)}=2\)

OpenStudy (anonymous):

oh well lets finish this one first coordinate of the vertex is \(2\) second coordinate of the vertex is what you get when you replace \(x\) by \(2\) namely \[-2(2)^2+8\times 2-5=3\] vertex is \[(2,3)\]

OpenStudy (anonymous):

can you fan me so i can message you

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Find the vertex of the parabola whose equation is \(y = x^2 + 6x + 2\) what is \(-\frac{b}{2a}\) in this case? hint: \(a=1,b=6\)

OpenStudy (anonymous):

you got this? or you still need help with it?

OpenStudy (anonymous):

hm, so c is 2?

OpenStudy (anonymous):

yes, \(c=2\) but you don't use \(c\) to compute the first coordinate of the vertex it is \[-\frac{b}{2a}\]

OpenStudy (anonymous):

for question one it is \[-\frac{6}{2\times (1)}=-3\]

OpenStudy (anonymous):

so 6/2 = 3

OpenStudy (anonymous):

don't forget the minus sign

OpenStudy (anonymous):

oh and the negative

OpenStudy (anonymous):

what about the other one?

OpenStudy (anonymous):

right now you have to find the second coordinate of the vertex \[y = x^2 + 6x + 2. \] and replace \(x\) by \(-3\) using parentheses gives you \[y=(-3)^2+6\times (-3)+2=9-18+2=-9+2=-7\]

OpenStudy (anonymous):

that makes the vertex \((-3,-7)\)

OpenStudy (anonymous):

what is the next one we need to do?

OpenStudy (anonymous):

wow, youre so quick. the next one is What is the line of symmetry for the parabola whose equation is y = 3x^2 + 24x - 1?

OpenStudy (anonymous):

i think the answer is -8 but im not sure

OpenStudy (anonymous):

once again it is \(-\frac{b}{2a}\) which in this case is \[-\frac{24}{2\times (3)}=-\frac{24}{6}\]

OpenStudy (anonymous):

so the answer is -4

OpenStudy (anonymous):

so no, it is not \(-8\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

since it is a line, the answer is actually \(x=-4\)

OpenStudy (anonymous):

what next?

OpenStudy (anonymous):

What is the line of symmetry for the parabola whose equation is y = x2 - 12x + 7?

OpenStudy (anonymous):

this one is for you to try first again it is \(-\frac{b}{2a}\) this time with \(a=1,b=-12\)

OpenStudy (anonymous):

so its 12/2

OpenStudy (anonymous):

yes better known as \(6\) but yes also don't forget it is \(x=6\) the vertical line

OpenStudy (anonymous):

that it, or is there more?

OpenStudy (anonymous):

ill make a new post

OpenStudy (anonymous):

k